OpenStudy (anonymous):

5^(3x) = 25^(x+1)? I got x=1 but I don't think that's right..

OpenStudy (anonymous):

It's not, try to check it again.

OpenStudy (anonymous):

I got 5^3x = 5^2x=1 then I did: 3x=2x+1 then I got x=1 :(

OpenStudy (anonymous):

2(x+1) is not equal to 2x +1 ..

OpenStudy (anonymous):

Oh I see lol I know the answer. You would want to use logarithms

OpenStudy (anonymous):

\[\eqalign{ &5^{3x}=25^{x+1} \\ &5^{3x}=(5^2)^{x+1}\\ &5^{3x}=5^{2x+2} \\ &3x=2x+2 \\ &x=2 }\]

OpenStudy (anonymous):

@KeithAfasCalcLover you are giving answers right away :(

OpenStudy (anonymous):

i'm trying to help him solve his problem ..

OpenStudy (anonymous):

oh yes I didn't multiply it out correctly! thanks @Jaseem18

OpenStudy (anonymous):

you are welcome @shaa

OpenStudy (anonymous):

don't worry I understand where I went wrong now, im a girl* haha

OpenStudy (anonymous):

The you just gave the answer though...He had the right method up until the point and you corrected him...That's all...Besides. the fourth step was the issue and I skipped the simplifying of that. There is not really any answer giving away.